OpenStudy (ninjamaster115):

Which ordered pair in the form (a, b), is a solution of this equation: 7a-5b=28 A. (-3, -2) B. (4, 0) C. (0, 4) D. (-2, -3)

OpenStudy (anonymous223):

(4,0) a=4 b=0 7-4-5x0=28 28-0=28 28=28

OpenStudy (anonymous223):

sorry i ment 7x4-5x0=28

OpenStudy (ninjamaster115):

its ok